6-3x+12=0 is the problem. I really need help understanding how to break down algebra equations
Answer by oscargut(2103)

6-3x+12=0
6+12-3x = 0
18-3x =0
18 = 3x
3x = 18
x = 18/3
x = 6
You can ask me more at: mthman@gmail.com
Thanks